OpenStudy (anonymous):

For what value(s) of k will the equation x^2 -12x + k = 0 have just one real root?

OpenStudy (lgbasallote):

if it has only one real root that means b^2 - 4ac = 0 so substitute a = 1 b = -12 c = k so (-12)^2 - 4(1)(k) = 0 144 - 4k = 0 add 4k to both sides 144 = 4k divide both sides by 4 36 = k does that help?

OpenStudy (anonymous):

Yes very much. Thankyou. And if there were more than one value it would be the same equation correct?

OpenStudy (lgbasallote):

if it has two real roots then you use b^2 - 4ac > 0 if it doesnt have any real root you use b^2 - 4ac <0

OpenStudy (anonymous):

Oh ok! Thank you so much! I appreciate it!

OpenStudy (lgbasallote):

welcome